OpenStudy (anonymous):

I have a problem understanding converting the integral of x^2 exp(ax^2) into polar form.

OpenStudy (anonymous):

Could anyone explain it?

OpenStudy (anonymous):

As much as I understand:\[I=\int\limits_{-\infty}^{\infty}x^2 e^{ax^2}dx\] \[I^2=\int\limits_{-\infty}^{\infty}\int\limits_{-\infty}^{\infty}x^2 e^{a(x^2+y^2)}dxdy\] Fro, here I fail to understand.

OpenStudy (anonymous):

*m

OpenStudy (anonymous):

Or even where the first x^2 is just x^1, for simplicity.

OpenStudy (anonymous):

that integral u wrote up there diverges ...

OpenStudy (anonymous):

Or from 0 to infinity, my problem here isn't the exact integral, but how you convert it into polar, and why.

OpenStudy (anonymous):

the only integral in that form i know is\[I=\int\limits_{-\infty}^{\infty} e^{-ax^2}dx\]with \(a>0\)

OpenStudy (anonymous):

Fine, then- just how do you convert it into polar (I can't remember the exact problem, sorry)?

OpenStudy (anonymous):

np henpen\[I^2=\int\limits_{-\infty}^{\infty} \int\limits_{-\infty}^{\infty} e^{-a(x^2+y^2)} \text{d}x \ \text{d}y\]so i think no problem till here

OpenStudy (anonymous):

Yes. I know that \[y^2+x^2=r^2=q\], (but whether r or q we will use here I do not know). Do we just substitute?

OpenStudy (anonymous):

yeah integrating \(f(x,y)\) in cartesian with dxdy equals to integrating \(f(r,\theta)\) in polar with rdrdtheta and of course we must change integrals bounds

OpenStudy (anonymous):

So it's not a case of robotically substituting and inserting the new values for dx, dy, y and x?

OpenStudy (anonymous):

this is like a doing sub but with 2 variable

OpenStudy (anonymous):

I get the intuition, but I still feel shaky. What is the proof of this (or, what f(x,y) is equal to an g(r, theta))?

OpenStudy (anonymous):

I mean, with substituting, r=x^2+y^2, dx=dr/2x, dy=dr/2y. Where does theta fit in to this?

OpenStudy (anonymous):

my friend herp-derp will clarify us

OpenStudy (anonymous):

Very kind of you, herp-derp.

OpenStudy (anonymous):

\[\text{If}~~I=\int_0^\infty x^2 e^{-ax^2}\text{d}x\]\[\text{then}~~I^2=\left(\int_0^\infty x^2 e^{-ax^2}\text{d}x\right)\left(\int_0^\infty y^2 e^{-ay^2}\text{d}y\right)=\int_0^\infty\int_0^\infty x^2y^2e^{-a(x^2+y^2)}\,\text{d}x\]You wrote it as \(\int_0^\infty\int_0^\infty x^2 e^{-a(x^2+y^2)}\text{d}x\,\text{d}y\) . Just wanted to clarify!

OpenStudy (anonymous):

Sorry, not really answering your question...

OpenStudy (anonymous):

Is that a dy missing from the end of your long equation?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I doubt it's a straight substitution, but what's the general rule for where you can/can't polarly substitute (and how)?

OpenStudy (anonymous):

It's here http://en.wikipedia.org/wiki/Polar_coordinate_system#Generalization , but I don't understand all of it.

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/CalcIII/DIPolarCoords.aspx depends on our integral...sometimes changing to polar will make the integral very easy to solve ...sometimes no... there is no general rule

OpenStudy (anonymous):

There must be a rule for conversion. I'll read the link.

OpenStudy (anonymous):

for the particular case of\[I^2=\int\limits_{-\infty}^{\infty} \int\limits_{-\infty}^{\infty} e^{-a(x^2+y^2)} \text{d}x \ \text{d}y\]thee integral becomes to\[I^2=\int\limits_{0}^{2\pi} \int\limits_{0}^{\infty} e^{-ar^2} r\ \text{d}r \ \text{d}\theta\]which is very easy to solve because of the 'r' that changing to polar provides

OpenStudy (anonymous):

People - get awareness finally that the x^2 factor at the start of the integrand equals\[(r* Cos \theta)^2\]

OpenStudy (anonymous):

Soo overall u obtain \[\int\limits_{-\infty}^{\infty}\int\limits_{-\infty}^{\infty}r^2 Cos^2 \theta \exp^{-\alpha r^2} r dr d \theta\]

OpenStudy (anonymous):

However Mr. H.D. here gave a smart and unobvious path to the solution itself

OpenStudy (anonymous):

However his path MUST cite atheorem allowing the double integral to equal a product of usual integrals

OpenStudy (anonymous):

henpen, for your case it is best to use a technique called "differentiation through the integral", because a change to polar coordinates will just make this integral messy. So notice that \(\displaystyle \int_0^\infty x^2e^{-ax^2}\,\text{d}x\) is equivalent to \(\displaystyle\int_0^\infty -\frac{\partial}{\partial a}\left[e^{-ax^2}\right]\,\text{d}x\). Using our knowledge of differentiation through the integral, we know that the last expression is equal to \(\displaystyle -\frac{\partial}{\partial a}\int_0^\infty e^{-ax^2}\,\text{d}x\). So what we need is that integral. Well, I will not explain it, but using a change to polar coordinates we can show that \(\displaystyle\int_0^\infty e^{-ax^2}\,\text{d}x=\sqrt{\frac{\pi}{4a}}\). Thus, \(\displaystyle -\frac{\partial}{\partial a}\int_0^\infty e^{-ax^2}\,\text{d}x=\frac{1}{4}\sqrt{\pi}a^{-3/2}\), which is your answer.

OpenStudy (anonymous):

Hey You knew ! You knew! I am betting you read this BRILLIANT move somwh. !

OpenStudy (anonymous):

Oh, I made a mistake. The derivative operator on the outside of the integral is not a partial derivative. It should read \(\displaystyle \frac{\text{d}}{\text{d}a}\int_0^\infty e^{-ax^2}\,\text{d}x\). Not that it matters.

OpenStudy (anonymous):

@Mikael Tis true, I did not come up with this method by myself. \(:(\)

OpenStudy (anonymous):

Genius method!

OpenStudy (anonymous):

@henpen Do you understand differentiation through the integral?

OpenStudy (anonymous):

First - give him another medal deserves so Second Her Derp - u must know som probability distrib or physics

OpenStudy (anonymous):

Yes, it's really elegant!

OpenStudy (anonymous):

Integrals of the form:\[\int_a^b \frac{\partial^{\,n}}{\partial^{\,n} x}\Big[f(x,\xi)\Big]\,\text{d}\xi=\frac{\text{d}^n}{\text{d}^nx}\int_a^b f(x,\xi)\,\text{d}\xi\]

OpenStudy (anonymous):

Basically, my original question is answered in that dA=rdrdtheta in general. Thanks for the link, mukshala!

OpenStudy (anonymous):

Why does it switch from partial to 'normal'?

OpenStudy (anonymous):

anytime henpen

OpenStudy (anonymous):

And if it's from a to b (i.e. definite), surely the right hand side is =0

OpenStudy (anonymous):

\(\displaystyle\int_a^b f(x,\xi)\,\text{d}\xi\) is \(\underline{\mathbf{not}}\) a function of \(\xi\). However, it \(\underline{\mathbf{is}}\) a function of \(x\).

OpenStudy (anonymous):

And sorry, mukshala, I think you had actually explained this to me before, but I forgot...

OpenStudy (anonymous):

Or- can the bounds be written as\[\xi(a)and \xi(b)\]? I think I get it.

OpenStudy (anonymous):

Imagine a 3-dimensional surface defined by \(z=f(x,y)\). Imagine a strip on the x-y plane from y=a to y=b. Now imagine a cross-section of this surface parallel to the x-z plane. This cross-section forms a curve. Finally, imagine finding the area under that curve. As you move the cross-section, the curve changes, and so does its area. In such a way, the area of the curve depends on the x-coordinate of the cross section. It doesn't depend on any y or z-coordinate.

OpenStudy (anonymous):

I did an example, and now I undertand that it works. You are an artful explainer, thank you a lot.

OpenStudy (anonymous):

I went to an art school for 3 years. (It was hell)